# EVE: Towards End-to-End Video Subtitle Extraction with Vision-Language Models

### v1

@@ 3.1 Architecture
Input : Predicted subtitles $\mathcal{S}$={$S_{0}$, $S_{1}$, …, $S_{T-1}$}, where ; OCR results of all frame in the input video $\mathcal{O}$={$O_{0}$, $O_{1}$, …., $O_{L-1}$}, where $O_{l}$ = $\text{o}_{0}$, $\text{o}_{1}$,…, $\text{o}_{K_{l}-1}$}; Similarity threshold $\mathbf{sim}$, Raw FPS of the input video $\mathbf{f}$, Subtile matching range $\mathbf{r}$
Output : Subtitles in SRT format $\mathcal{S}^{\prime}$={$S^{\prime}_{0}$, $S^{\prime}_{1}$, …, $S^{\prime}_{T-1}$}, $S^{\prime}_{t}$=[$b^{\prime}_{t}$, $e^{\prime}_{t}$, $s^{\prime}_{t}$]
@@
Algorithm 1 Convert EVE predictions into subtitles in SRT format.
@@
Large Language Model. In the past few years, numerous foundation large language models are open-sourced, such as Qwen [40] and LLaMa [29]. Since the proposed EVE aims at solving bilingual (i.e., English and Chinese) video subtitle extraction, we select Qwen as the large language model in EVE. In addition, taking into account the time efficiency in practical applications, we build EVE-7B and EVE-0.5B based on Qwen2-7B and Qwen2-0.5B, respectively. Detail performance comparison is shown in Sec. 5.
3.2 Training
@@
The training process of the proposed EVE is divided into two stages: pre-training stage and supervised fine-tuning stage. We compare these two different stages from the following three aspects:
@@
  • Training Details. The pre-training stage primarily aims at enhancing the text perception capability of EVE, enabling it to predict all text tracklets in video frames. As shown in Figure 3, at the pre-training stage, the large language model is frozen while the vision encoder and the proposed adapter InterleavedVT are optimized. After the pre-training stage, the vision encoder is well trained to perceive texts. Thus, at the supervised fine-tuning stage, the vision encoder is frozen while the adapter InterleavedVT and the language model are optimized to learn discriminate whether texts in the input video are subtitles.
@@
  • Input. Both pre-training and supervised fine-tuning stages share the same input template. In the template, we additionally prepend the corresponding frame index to the visual tokens to improve the temporal perception ability of our model. The limit for the number of input frames is set to 30 during pre-training while it is set to 120 during supervised fine-tuning.
@@
  • Prediction. For prediction, the two stages also share the same template, where the start/end frame index and texts are included. Differently, the pre-training stage learns to recognize all text tracklets in videos while the supervised fine-tuning stage only considers those subtitle tracklets.
@@
Figure 3: Training details of EVE. The LLM is frozen during pre-training while the vision encoder is forzen during supervised fine-tuning.
3.3 Inference
@@
Due to the limitation of EVE in input token length, inference cannot be completed at high frame rates. Therefore, the proposed EVE samples the video at 2 frames per second (FPS) for model input. Although the predicted frame index can directly correspond to a precise timestamp (i.e., hour, minute, and second), it does not represent the precise time at which each subtitle appears and disappears in the video. To output more accurate subtitle content, such as in SRT format (examples are provided in the supplementary materials), we propose a post-processing method based on OCR results at full frame rate. The algorithm of post-processing is described in Algorithm 1. Firstly, we use an OCR engine to detect and recognize text in images at full frame rate. Then, we map the frame indices identified by EVE to rough timestamps at full frame rate. Finally, by calculating the edit distance between predicted texts and those within adjacent frames, the precise subtitle timestamp can be obtained.
@@
Table 1: The number of samples in different languages. Both short videos and movie videos are mainly in Chinese.
4 Data Construction
@@
In this paper, we propose a large dataset, called ViSa, to benchmark the video subtitle extraction task. The raw videos are collect from two datasets: CNVid [9] and MovieChat-1K [27]. CNVid is the largest public video-text dataset in Chinese, which contains 3.5M short videos collect from Douyin  [^5] https://www.douyin.com/. MovieChat-1K contains 1K long movie videos with bilingual subtitles.
@@
Based on these two datasets, we construct ViSa for video subtitle extraction. For CNVid, we perform the following three steps to reserve the ideal videos: (1) First, we filter out videos with a duration shorter than 10 seconds or longer than 120 seconds. The reason for this operation is that short videos (duration$<$10s) contain a lot of noise while the number of frames in long videos (duration$>$120s) exceeds the model’s input limits. (2) Then, we utilize a in-house video OCR tools to predict the text tracklets in the reserved videos. (3) Finally, videos with less than five tracklets will be filtered out, as these videos are unlikely to contain subtitles. Due to the fact that the movie videos in MovieChat-1K are mostly longer than 30 minutes, we clip each movie video into multiple short videos with a random duration (from 15s to 60s). By processing the above two datasets, we can obtain 2.5M short videos to build the video subtitle extraction dataset ViSa.
@@
Figure 4: The statistics information of ViSa. The proportion of videos without subtitles in movie videos is higher that in short videos. Since we clip movie videos based on a uniformly sampled duration in the range of 15--60 seconds, the duration distribution of movie videos is more balanced compared to short videos.
@@
Specifically, ViSa divides the collected videos into three parts: pre-training, supervised fine-tuning, and test datasets. Some statistics of them are described in the following. More details are shown in the supplementary materials.
@@
Pre-training dataset. The pre-training dataset contains 2.35M videos, all of which are collected from CNVid. The videos in the pre-training dataset are annotated with text tracking and text recognition since the goal of pre-training is to equip the model with the text perception capability. The statistics of pre-training dataset are shown in the supplementary materials. The majority of pre-training videos are typically shorter than 60 seconds in length. The number of text tracklets in videos mostly ranges between 12 and 30. In a total of 2.5M pre-training videos, 1.8M videos include text tracklets while 0.7M videos contain no texts.
@@
Supervised Fine-tuning (SFT) dataset. The videos for supervised fine-tuning originate from both CNVid and MovieChat-1K. Concretely, the SFT dataset includes 210K Chinese videos from CNVid and 20K bilingual videos from MovideChat-1K. Each video in the SFT dataset is annotated with subtitle text tracking and corresponding results. As a data augmentation strategy to expand the SFT dataset, we annotate each video with both 1-FPS and 2-FPS sampling rate. The statistics of SFT dataset are shown in Fig. 4. As shown in Tab. 1, among 230K SFT videos, 30K videos contains no texts.
@@
Test dataset. To evaluate the performance of EVE on both short videos and movies, we collect 857 videos from CNVid and 862 videos from MovieChat-1K to build the test dataset. The annotations of this dataset are the same as that of the SFT dataset and they are manually annotated. Similar to the SFT dataset, both positive samples (i.e., videos containing texts) and negative samples (i.e., videos containing no texts) are also included in the test dataset.
@@
Open-sourced Tools VideOCR [4] 0.348 72.72 52.63 0.686 6.28 57.12
VideOCR [4] + Subregion 0.659 52.92 43.81 0.702 5.24 66.35
PyVideoTrans [3] 0.477 46.03 92.34 0.639 88.35 52.51
Cloud Service OpenVision [2] 0.771 21.85 42.85 0.718 6.57 18.61
GhostCut [1] 0.655 54.78 46.00 0.659 7.26 24.43
Open-sourced LVLMs Qwen2-VL-2B [31] 0.068 - - 0.070 - -
Qwen2-VL-7B [31] 0.123 - - 0.227 - -
MiniCPM-V-2.6 [42] 0.043 - - 0.047 - -
@@
Table 2: Performance comparison with existing three types of methods. Since the open-sourced LVLMs are trained with no video subtitle extraction datasets, they cannot achieve ideal performance on this task.
5 Experiments
5.1 Implementation Details
The proposed method is implemented with PyTorch. All experiments are conducted on $24$ A100 GPUs. The proposed EVE is trained using the AdamW optimizer [20] where the learning rate is set to $5\times 10^{-5}$. The input frames are resized into $640\times 640$. The batch size is set to 48 for pre-training and 16 for supervised fine-tuning. The number of queries $K$ in the textual region compressor is set to 10, and the number of visual tokens is set to 16. More relevant experimental results are shown in the supplementary materials. The vision encoder is initialized with the pre-trained parameters of Donut [11], and the large language model loads the pre-trained models of Qwen2-7B or Qwen2-0.5B.
5.2 Metrics
In this paper, we adopt two types of metrics to evaluate the performance of models: Normalized Edit Distance (NED) [44] and SubER [33]. NED is defined as:
@@ 5.2 Metrics
where ED($\cdot$) denotes the edit distance function. The function MaxL(A, B) returns the maximum length between strings A and B. $S$ represents the number of samples. The NED metric is utilized to calculate the accuracy of predicting subtitle texts. However, it may hardly evaluate the accuracy of timestamp predictions. Therefore, we additionally adopt SubER [33] to evaluate the performance of models. SubER is defined as:
@@
where “#” means “number of”. word edits are insertions, deletions and substitutions of words; break edits are insertions, deletions and substitutions of line breaks; shifts are movements of one or more adjacent hypothesis tokens to a position of a matching phrase in the reference; reference words/reference breaks is the ground-truth words/breaks. More details about SubER can refer to [33].
5.3 Experimental Results
Evaluating Existing Methods. Although the video subtitle extraction task has been explored before, there is still a lack of comprehensive dataset to benchmark this task. In this paper, we propose a video subtitle extraction dataset containing pre-training, supervised fine-tuning and test datasets. We have evaluated three categories of video subtitle extraction methods on the proposed test dataset. Among the open-sourced tools, PyVideoTrans based on ASR achieves relatively inferior performance on negative samples compared with VideOCR. The reason for this results is that the audio in the negative sample will also be translated into sentences, which are regarded as subtitles of the input video. We have also evaluated two cloud services that can perform video subtitle extraction. Ghost cannot perform well on short videos containing a lot of background texts since it tends to mistake background texts as subtitles. In contrast, OpenVision can achieve relatively ideal performance on both short videos and movie videos. Two open-source multi-modal large models (Qwen2-VL [31] and MiniCPM-V-2.6 [42]) perform poorly on the video subtitle extraction task because they are not trained on corresponding datasets. They prefer to performing video captioning rather than extracting subtitle texts, thus resulting in subpar performance. In addition, the experimental results demonstrate that Qwen2-VL’s video OCR capability is better than MiniCPM-V-2.6. More visualizations are shown in supplementary materials.
Comparison with Existing Methods. In this paper, we establish two models with different LLMs. To achieve better performance in video subtitle extraction, we utilize Qwen2-7B to construct EVE-7B. Through the experimental results in Tab. 2, EVE-7B achieves the best performance compared with three types of video subtitle extraction methods. For example, on the short video dataset, EVE-7B outperforms OpenVision by 11.1% in the metric of NED. We have visualized an example in Fig 5, and the result shows that the proposed EVE can well predict the subtitles without missing and repetition problem. More visualizations are shown in the supplementary materials. We have also conducted experiments to compare with two typical token compression methods (T-Selector [30] and ToMe [5]). The results are shown in the supplementary materials.
Considering that the EVE-7B will cost more computing resources in practical applications, we establish a light-weighted model EVE-0.5B with Qwen2-0.5B. The experimental results in Tab. 2 demonstrate that despite using the same dataset for training, EVE-0.5B still cannot achieve the same performance as EVE-7B. However, benefiting from the proposed InterleavedVT and corresponding dataset ViSa, it can still surpass most of existing methods.
Figure 5: An example of subtitle extraction for different methods. The ASR based method PyVideoTrans tends to miss some audio-irrelative texts; Qwen2-VL performs video captioning rather than extracting subtitles; OpenVision may be based on text tracking, resulting repetition problem when text tracklets are wrongly predicted. Gray and Red indicate the missing and repetition prediction problem, respectively. Green represents the format of the answer is incorrect.
VC TRC SP Short Video Movie

### v2

@@ III-C3 Feature Alignment and Fusion
Input : Predicted subtitles $\mathcal{S}$$S_{0}$, $S_{1}$, …, $S_{T-1}$}; OCR results of all frame in the input video $\mathcal{O}$$O_{0}$, $O_{1}$, …., $O_{L-1}$}, where $O_{l}$ = $\text{o}_{0}$, $\text{o}_{1}$,…, $\text{o}_{K_{l}-1}$}; Similarity threshold $\mathbf{sim}$; Raw FPS of the input video $\mathbf{f}$; Subtile matching range $\mathbf{r}$.
Output : Subtitles in SRT format $\mathcal{S}^{\prime}$$S^{\prime}_{0}$, $S^{\prime}_{1}$, …, $S^{\prime}_{T-1}$}, $S^{\prime}_{t}$$b^{\prime}_{t}$, $e^{\prime}_{t}$, $s^{\prime}_{t}$]
1ex
@@
Algorithm 1 Converting EVE predictions into subtitles with SRT format.
III-D Large Language Model
@@
In the past few years, numerous foundation large language models are open-sourced, such as Qwen2 [yang2024qwen2] and LLaMa [touvron2023llama]. Since the proposed EVE aims at solving bilingual (i.e., English and Chinese) video subtitle extraction, we select Qwen2 as the large language model of EVE. In addition, taking into account the time efficiency in practical applications, we build EVE-7B and EVE-0.5B based on Qwen2-7B and Qwen2-0.5B, respectively. Detail performance comparison is shown in Section V.
III-E Training
@@
The training process of the proposed EVE consists of two stages: a pre-training stage and a supervised fine-tuning stage. We describe these two stages in detail from the following three aspects:
@@
  • Training Details. The pre-training stage aims to enhance the text perception capability of EVE by enabling it to predict all text tracklets appearing in video frames. As shown in Figure 3, during this stage, the large language model is frozen, while the vision encoder and the proposed S3 module are jointly optimized. After pre-training, the vision encoder becomes proficient at perceiving textual regions. Consequently, in the supervised fine-tuning stage, the vision encoder is frozen, and only the S3 module together with the large language model are trained to distinguish whether the detected texts correspond to subtitles.
@@
  • Input. Both stages share the same input template. To improve temporal perception, we prepend the corresponding frame index to each frame’s visual tokens. The number of input frames is limited to 30 during pre-training and 120 during supervised fine-tuning.
@@
  • Prediction. The prediction template is also identical across the two stages, containing the start frame index, end frame index, and texts. The difference lies in their objectives: the pre-training stage learns to recognize all text tracklets in videos, whereas the supervised fine-tuning stage focuses solely on those corresponding to subtitles.
@@
Figure 3: Training details of EVE. The LLM is frozen during pre-training while the vision encoder is frozen during supervised fine-tuning.
III-F Inference
III-F1 Output Format
We use the SRT files as the model’s output, as they are one of the most popular subtitle file formats for video content. An SRT file stores sequential subtitle texts with corresponding timestamps but no audio or video data, resulting in a compact file size. In addition, the SRT file needs to specify the line number, subtitle content, when the text should appear on the video, and when it should disappear to make room for the next line, as shown in Figure 4.
III-F2 Frame Sampling and Post-processing
@@
Due to the limitation of EVE in input token length, inference cannot be completed at high frame rates. Therefore, the proposed EVE samples the video at 2 frames per second (FPS) for model input. Although the predicted frame index can directly correspond to a precise timestamp (i.e., hour, minute, and second), it does not represent the precise time at which each subtitle appears and disappears in the video. To output more accurate subtitle content, such as in SRT format, we propose a post-processing method based on OCR results at full frame rate. The algorithm of post-processing is described in Algorithm 1. Firstly, we use an OCR engine to detect and recognize text in images at full frame rate. Then, we map the frame indices identified by EVE to rough timestamps at full frame rate. Finally, by calculating the edit distance between predicted texts and those within adjacent frames, the precise subtitle timestamp can be obtained.
Example of a bilingual SRT subtitle file 1 00:00:01,739 $\rightarrow$ 00:00:02,826 那么我是谁? Then who am I? 2 00:00:03,521 $\rightarrow$ 00:00:04,913 你就是那个节目的明星 You’re the star. 5 00:00:04,540 $\rightarrow$ 00:00:05,826 所以才有那么多人看你 That’s what made you so good to watch.
Figure 4: An example of a bilingual SRT subtitle file.
@@
TABLE I: The number of samples in different languages. Both short videos and movie videos are mainly in Chinese.
@@
IV Dataset
Although video subtitle extraction has been studied for years, existing datasets for this task remain extremely limited. To this end, we present the ViSa dataset, the first large-scale dataset for video subtitle extraction, providing training material and a evaluation benchmark dataset for this task. In accordance with the previously mentioned training objectives, we divide ViSa into three components: the pre-training dataset, the supervised fine-tuning dataset, and the test dataset.
IV-A Comparison With Existing Dataset
Elshahaby et al. [elshahaby2022end] introduced the FiViD dataset for video subtitle extraction task, while Yan et al. [yan2020end] trained and evaluated their model on publicly available image-based OCR datasets, including MJ Synth [jaderberg2014synthetic], ICDAR03 [1227749], ICDAR13 [6628859], and SVT [Wang20111457]. In contrast, Mathur et al. [mathur2015generating] relied on the audio-only VoxForge [voxforge] corpus to explore speech-to-text subtitle generation, a direction that falls outside the scope of the visual subtitle extraction task studied in this work. As shown in Table  II, the aforementioned datasets are typically limited in scale and lack timestamp annotations. Moreover, most of them composed of images rather than video. Consequently, these datasets are insufficient to support modern LVLM-based models that require large-scale, high-quality supervision, especially for fine-grained temporal localization.
Meanwhile, recent large video datasets such as CNVid [gan2023cnvid] and MovieChat-1K [song2024moviechat] provide massive video–text pairs, yet they are not designed for subtitle extraction. They contain video captions or movie subtitles that (1) do not align with frame-level timestamps, (2) are not guaranteed to reflect on-screen subtitle content, and (3) lack start–end boundaries required by this task. Thus, they cannot be directly used to train or evaluate a subtitle extraction model.
To fill this gap, we construct ViSa, the first large-scale dataset specifically designed for timestamped video subtitle extraction. ViSa contains bilingual (Chinese and English) subtitle texts paired with both coarse-grained and fine-grained timestamps, enabling scalable pre-training, supervised fine-tuning, and rigorous evaluation of video subtitle extraction systems. Table II summarizes the differences between ViSa and prior subtitle extraction datasets.
TABLE II: Comparison of ViSa with datasets used in previous works.
Dataset
Data
Type
Scale
On-screen Text Annot.
Time-
stamp
FiViD [elshahaby2022end] Image $\approx$74k Yes No
MJ Synth [jaderberg2014synthetic] Synthetic 9M Yes No
ICDAR03 [1227749] Image 860 Yes No
ICDAR13 [6628859] Image 857 Yes No
SVT [Wang20111457] Image 647 Yes No
CNVid [gan2023cnvid] Video 3.5M No No
MovieChat-1K [song2024moviechat] Video 1k No No
ViSa (ours) Video 2.5M Yes Yes
IV-B Data Construction
@@
The raw videos of ViSa are collected from two datasets: CNVid [gan2023cnvid] and MovieChat-1K [song2024moviechat]. CNVid is the largest public video-text dataset in Chinese, which contains 3.5M short videos collect from Douyin  [^6] https://www.douyin.com/. MovieChat-1K contains 1K long movie videos with bilingual subtitles.
@@
Based on these two datasets, we construct ViSa for video subtitle extraction. For CNVid, we perform the following three steps to reserve the ideal videos: (1) First, we filter out videos with a duration shorter than 10 seconds or longer than 120 seconds. The reason for this operation is that short videos (duration<<10s) contain a lot of noise while the number of frames in long videos (duration>>120s) exceeds the model’s input limits. (2) Then, we utilize a in-house video OCR tools to predict the text tracklets in the reserved videos. (3) Finally, videos with less than five tracklets will be filtered out, as these videos are unlikely to contain subtitles. Due to the fact that the movie videos in MovieChat-1K are mostly longer than 30 minutes, we clip each movie video into multiple short videos with a random duration (from 15s to 60s). By processing the above two datasets, we can obtain 2.5M short videos to build the video subtitle extraction dataset ViSa.
@@
Figure 5: The statistics information of ViSa. The proportion of videos without subtitles in movie videos is higher that in short videos. Since we clip movie videos based on a uniformly sampled duration in the range of 15-60 seconds, the duration distribution of movie videos is more balanced compared to short videos.
@@
Specifically, ViSa divides the collected videos into three parts: pre-training, supervised fine-tuning, and test datasets. Some statistics of them are described below.
IV-C Pre-training dataset.
@@
The pre-training dataset contains 2.35M videos, all of which are collected from CNVid. The videos in the pre-training dataset are annotated with text tracking and text recognition since the goal of pre-training is to equip the model with the text perception capability. Therefore, this annotation is coarse-grained and low-cost, considering the large scale of this portion of the data. The majority of pre-training videos are typically shorter than 60 seconds in length. The number of text tracklets in videos mostly ranges between 12 and 30. In a total of 2.35M pre-training videos, 1.8M videos include text tracklets while 0.7M videos contain no texts.
IV-D Supervised Fine-Tuning dataset.
@@
The videos for Supervised Fine-Tuning (SFT) originate from both CNVid and MovieChat-1K. Concretely, the SFT dataset includes 210K Chinese videos from CNVid and 20K bilingual videos from MovieChat-1K. Each video in the SFT dataset is annotated with subtitle text tracking and corresponding results.
Specifically, we use the annotation tool arctime  [^7] https://arctime.org to perform automated annotation for each video, followed by manual review and correction to ensure the subtitles are accurate and the timestamp error is less than 0.5 seconds. As a data augmentation strategy to expand the SFT dataset, we annotate each video with both 1-FPS and 2-FPS sampling rate. The statistics of SFT dataset are shown in Figure 5. As shown in Table I, among 230K SFT videos, 30K videos does not contain text.
IV-E Test dataset.
@@
To evaluate the performance of EVE on both short videos and movies, we collect 857 videos from CNVid and 862 videos from MovieChat-1K to build the test dataset. The annotation for this portion of the data is fine-grained and manually generated to enable objective model evaluation. Similar to the SFT dataset, both positive samples (i.e., videos containing texts) and negative samples (i.e., videos containing no texts) are also included in the test dataset.
TABLE III: Performance comparison with existing three types of methods.
@@
Open-sourced Tools VideOCR [videocr] 0.348 72.72 52.63 0.686 6.28 57.12
VideOCR [videocr] + Subregion 0.659 52.92 43.81 0.702 5.24 66.35
PyVideoTrans [pyvideotrans] 0.477 46.03 92.34 0.639 88.35 52.51
Cloud Service OpenVision [openvision] 0.771 21.85 42.85 0.718 6.57 18.61
GhostCut [ghost] 0.655 54.78 46.00 0.659 7.26 24.43
Open-sourced LVLMs Qwen2-VL-2B [wang2024qwen2] 0.068 - - 0.070 - -
Qwen2-VL-7B [wang2024qwen2] 0.123 - - 0.227 - -
MiniCPM-V-2.6 [yao2024minicpm] 0.043 - - 0.047 - -
@@
V Experiments
V-A Implementation Details
The proposed method is implemented with PyTorch. All experiments are conducted on $24$ A100 GPUs. The proposed EVE is trained using the AdamW optimizer [loshchilov2017decoupled] where the learning rate is set to $5\times 10^{-5}$. The input frames are resized into $640\times 640$. The batch size is set to 48 for pre-training and 16 for supervised fine-tuning. The shape of $\mathbf{F}^{v}_{i}$ is set to $\mathbb{R}^{4\times 4\times C}$. The number of queries $Q$ in the TSTQ is set to 10, and the number of visual tokens after SSCA is set to 16. The vision encoder is initialized with the pre-trained parameters of Donut [kim2021donut], and the large language model loads the pre-trained models of Qwen2-7B or Qwen2-0.5B.
V-B Metrics
In this paper, we adopt two types of metrics to evaluate the performance of models: Normalized Edit Distance (NED) [zhang2019icdar] and SubER [wilken2022suber]. NED is defined as:
@@ V-B Metrics
where ED($\cdot$) denotes the edit distance function. The function MaxL(A, B) returns the maximum length between strings A and B. $S$ represents the number of samples. The NED metric is utilized to calculate the accuracy of predicting subtitle texts. However, it may hardly evaluate the accuracy of timestamp predictions. Therefore, we additionally adopt SubER [wilken2022suber] to evaluate the performance of models. SubER is defined as:
@@
where “#” means “number of”. word edits are insertions, deletions and substitutions of words; break edits are insertions, deletions and substitutions of line breaks; shifts are movements of one or more adjacent hypothesis tokens to a position of a matching phrase in the reference; reference words/reference breaks is the ground-truth words/breaks. More details about SubER can refer to [wilken2022suber].
TABLE IV: Performance comparison with two typical token compression methods.
Method Short Video Movie
Positive Negative Positive Negative
NED SubER SubER NED SubER SubER
T-Selector [wang2025elysium] 0.801 20.32 23.84 0.731 8.72 19.30
ToMe [bolyatoken] 0.714 28.35 25.62 0.705 10.79 15.59
EVE-0.5B 0.759 22.82 18.36 0.685 5.10 17.06
EVE-7B 0.882 15.16 13.92 0.808 1.64 10.15
V-C Experimental Results
V-C1 Evaluating Existing Methods
We evaluated three categories of video subtitle extraction methods on the proposed test dataset as shown in Table  III. Among the open-sourced tools, PyVideoTrans based on Automatic Speech Recognition (ASR) achieves relatively inferior performance on negative samples compared with VideOCR. The reason for these results is that ASR-based methods inherently treat spoken audio as potential subtitles, causing false positives in negative samples where no on-screen text exists. We also evaluated two cloud services that can perform video subtitle extraction. GhostCut cannot perform well on short videos containing a lot of background texts since it tends to mistake background texts as subtitles. In contrast, OpenVision can achieve relatively ideal performance on both short videos and movie videos. Two open-source multi-modal large models (Qwen2-VL [wang2024qwen2] and MiniCPM-V-2.6 [yao2024minicpm]) perform poorly on the video subtitle extraction task because they are not trained on corresponding datasets. They prefer to performing video captioning rather than extracting subtitle texts, thus resulting in subpar performance. In addition, the experimental results demonstrate that Qwen2-VL’s video OCR capability is better than MiniCPM-V-2.6.
TABLE V: Ablation studies on the proposed EVE (metric: SubER).
SSCA TSTQ SP Short Video Movie